OpenStudy (anonymous):

limx-->0 of (a-cosbx)/x^2 =8 find a and b without lhopital

OpenStudy (perl):

we know that sin x / x as x->0 =1 and

OpenStudy (perl):

and lim (1 - cos x ) / x = 0 as x->0

OpenStudy (perl):

I would use Lhopitals lim (-(-sin bx ) *b / (2x ) = 8

OpenStudy (anonymous):

i know using lhopital,but without it how can i start it

OpenStudy (perl):

let b = 4

OpenStudy (perl):

if b = 4 then limx-->0 of (a-cosbx)/x^2 =lim (-(-sin bx ) *b / (2x ) = lim sin(4x)*4 / (2x) = lim sin(4x)*4*2 / (2*2x) = lim sin(4x)/(4x) * 8 = 8 * lim sin(4x)/4x) = 8 *1 note that a = any constant . pick 0 if you want

OpenStudy (anonymous):

thank you

OpenStudy (perl):

see if you can generalize this problem

OpenStudy (perl):

limx-->0 of (a-cosbx)/x^2 =k find a,b

OpenStudy (zarkon):

I thought he wanted it without using L'Hospitals

OpenStudy (perl):

right

OpenStudy (perl):

so the question is still open

OpenStudy (zarkon):

clearly a=1 otherwise the limit would not exist then multiply by \(1+\cos(bx)\) to the top and bottom and simplify

OpenStudy (perl):

good point, if a equals zero that the limit diverges

OpenStudy (anonymous):

i was thinking about the squeeze thm, however the =8 is throwing me off. I wanted to say -1/x^2 <_ a-cosbx/x^2 <_ 1/x^2, but that does nothing

OpenStudy (perl):

n121 i would use zarkon's suggestion

OpenStudy (perl):

squeeze theorem probably won't be of use here

OpenStudy (perl):

Zarkon's suggestion makes a nice tidy simplification

OpenStudy (anonymous):

i will try

OpenStudy (perl):

also use the suggestion that a = 1, otherwise the limit diverges (do you see why?)

OpenStudy (anonymous):

yes